The sum of a number and 9 is subtracted from 60. The result is 10. Find the number. Type in your numerical answer only; do not type any words or letters with your answer
Answer by fractalier(6550)

60-(x+9)=10
60-x-9=10
51-x=10
-x = -41
x = 41